Calculus I, Fall 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 15

1. Section 8.3: Problems 3, 5, 13, 35, 37, 45, 47, 48, 54, 57.

2. Section 8.4: Problems 1, 3, 11, 13, 33, 39, 41, 49. Homework for next week (these are harder, but you can try first): Problems 15, 17, 19, 21, 23, 27, 29, 32, 35, 37, 43, 45.